OpenStudy (anonymous):

what should i do in this ... 2^8 × 8^2 × 4^-2 in the form 2n. ???

OpenStudy (raden):

try to get the bases of them be same. hint : 8 = 2^3 4 = 2^2

OpenStudy (skullpatrol):

Put it in the form 2n or 2^n ?

OpenStudy (anonymous):

2^n

OpenStudy (skullpatrol):

Bigger hint: $$2^8 × 8^2 × 4^{-2} $$ $$=2^8 × (2^3)^2 × (2^2)^{-2} $$

OpenStudy (anonymous):

there's another question....!! what is the formula of compund interest ? ;p

OpenStudy (raden):

super hint :P use these properties of exponents : (a^m)^n = a^(m * n) a^m * a^n = a^(m+n)

OpenStudy (skullpatrol):

formatted super hint: $$(a^m)^n = a^{m * n}$$and $$a^m * a^n = a^{m+n}$$

OpenStudy (gorv):

2^8 × 8^2 × 4^-2 2^8*(2^3)^2*(2^2)^-2 2^8*2^6*2^-4 as base of all the power is same that is 2 so the power will add up 2^(8+6-4) 2^10